Simplified equation of x^2 y^2 - 2x + 4y = 12
Answer by Boreal(15235)

I am editing this to read x^2+y^2-2x+4y=12
Rewrite as
x^2-2x+....+ y^2+4y+...=12
Now complete the square for each
x^2-2x+1+Y^2+4y+4=12+1+4, adding the 1 and the 4 to both sides.
(x-1)^2+(y+2)^2=17
This is a circle centered on (1,-2) with radius sqrt(17)